please help solve this equation?? I do not understand how to do this.
(5x+5)(10x-6)=0

Answer by ankor@dixie-net.com(22740)

(5x+5)(10x-6)=0
Make each factor equal to 0, and find x
5x + 5 = 0
subtract 5 from both sides
5x = -5
Divide both sides by 5
x = -1, is one solution
and
10x - 6 = 0
add 6 to both sides
10x = 6
divide both sides by 10
x = 6/10
x = .6 is the other solution